Please help me with solve by completing the square.
x^2+5^x-24=0

Answer by Alan3354(69443)

Please help me with solve by completing the square.
x^2+5^x-24=0 (I think you mean 5x)
-----------
x^2 + 5x = 24
c = (b/2)^2 = (5/2)^2
c = 6.25
Add c to both sides
x^2 + 5x + 6.25 = 30.25
(x + 2.5)^2 = 5.5^2
x + 2.5 = ± 5.5
x = -8
x = 3